OpenStudy (anonymous):

how to do this problem Line WX contains (−1, 2) and (4, 12) Line YZ contains points (5, 8) and (−2, −6). Lines WX and YZ ?

OpenStudy (anonymous):

we have to find the slope of these lines using the slope formula slope = (y2 - y1) / (x2 - x1) Line WX : (-1,2) x1 = -1 and y1 = 2 (4,12) x2 = 4 and y2 = 12 now sub into the slope formula slope = (12 - 2) / (4 - (-1) slope = 10/(4 + 1) slope = 10/5 which reduces to 1/2 so the slope of line WX = 1/2 Line YZ : (5,8) x1 = 5 and y1 = 8 (-2,-6) x2 = -2 and y2 = -6 now we sub into slope formula slope = (-6-8) / (-2-5) slope = -14/-7 slope = 1/2 so the slope in line YZ = 1/2 Line WX and line YZ are : (D) parallel because the slopes are the same.